Opened 9 years ago

Closed 5 years ago

#321 closed flight processing (fixed)

GB10/07, flight day 063/2010, Upavon

Reported by: benj Owned by: knpa
Priority: alpha 4 low Milestone: 2010 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by anch)

Data location: ~arsf/arsf_data/2010/flight_data/uk/GB10_07-2010_063_Upavon

Data arrived from ARSF via network transfer on 19th March 2010. Rinex data arrived separately via CD.

Scientific objective: Investigation of the effect of archaeological features on airborne hyperspectral and lidar imagery.

Priority: a4l

PI: Dr. Kate Welham

Sensors:

  • Eagle : Delivered 16/08/2010
  • Hawk : Delivered 20/12/2010
  • Leica LIDAR : Delivered 23/08/2010
  • Camera : Delivered 23/08/2010

Change History (32)

comment:1 Changed 9 years ago by benj

Note the nextmapdem.sh script generates a DEM for this flight containing tiles so90, su14 and su15. These are not adjacent and if you let the script do what it wants it will generate a DEM with a large blank area. This should be checked and corrected prior to processing this project.

comment:2 Changed 9 years ago by benj

  • Description modified (diff)
  • Priority changed from alpha 4 medium to alpha 4 low

Corrected priority to a4l.

comment:3 Changed 9 years ago by benj

  • Description modified (diff)

comment:4 Changed 9 years ago by benj

  • Milestone set to 2010 data processing completion

comment:5 Changed 9 years ago by pmlvis10

IPAS navigation for this project has been processed.

comment:6 Changed 9 years ago by ella

Looks like applanix navigation processing has been done. Copied to arsf workspace in Samba.

comment:7 Changed 9 years ago by anch

  • Owner set to anch
  • Status changed from new to accepted

Julian day supplied on the log sheet is 64 for some reason, but the date and julian day of the directory/ticket are correct.

comment:8 Changed 9 years ago by anch

SCT Offsets:

EAGLE
-2 : 0.02
-3 : 0.01
-4 : 0.04
-5 : 0.05
-6 : 0.01
-7 : 0.07
-8 : -0.03
-9 : 0.03
-10 : 0.05
-11 : 0.01
-12 : 0.02
-13 : 0.07
-14 : 0.02
-15 : -0.01
-16 : 0.05
-17 : -0.03

HAWK
-2 : 0.00
-3 : 0.01
-4 : -0.01
-5 : 0.00
-6 : 0.04
-7 : -0.01
-8 : 0.01
-9 : 0.00
-10 : 0.00
-11 : 0.01
-12 : -0.01
-13 : 0.02
-14 : -0.04
-15 : 0.08
-16 : -0.05
-17 : 1.1

Hyperspectral processing complete - submitted for delivery checking

comment:9 Changed 9 years ago by jb362

Beginning delivery check

comment:10 Changed 9 years ago by jb362

Delivery check complete

comment:11 Changed 9 years ago by anch

The project PI has requested the Eagle data be delivered before issues with Hawk have been resolved.

I have created a seperate delivery in folder 20100813 containing just Eagle data, which I will deliver ASAP.

comment:12 Changed 9 years ago by anch

  • Description modified (diff)

16/08/2010

Eagle data has been dispatched on disk via special delivery, Eagle data has also been rsynced, and I have emailed the PI to confirm dispatch of the data.

comment:13 Changed 9 years ago by anch

I've started to process the LiDAR for this project.

comment:14 Changed 9 years ago by emca

Camera has been processed and ready for delivery check at
~airborne/workspace/GB10_07-2010_063_Upavon/delivery/20100818/GB10_07

Removed images 1-19 as they did not cover the study area.

comment:15 Changed 9 years ago by jb362

delivery checking camera

comment:16 Changed 8 years ago by jb362

Delivery checking LiDAR

comment:17 Changed 8 years ago by jb362

LiDAR fine. Removed .svn directory from bin/pt_cloud_filter.

comment:18 Changed 8 years ago by jb362

Camera fine

comment:19 Changed 8 years ago by anch

rsynced LiDAR and camera to arsf_data

comment:20 Changed 8 years ago by anch

  • Description modified (diff)

23/08/2010

Sent HDD containing LiDAR and camera to PI.

HDD was number 70 which I have recorded.

Now we are just waiting on Hawk calibration

comment:21 Changed 8 years ago by emca

Delivery checked Hawk data - ready to deliver.

comment:22 Changed 8 years ago by anch

Hawk dispatched

comment:23 Changed 8 years ago by anch

  • Description modified (diff)

comment:24 Changed 8 years ago by anch

Hawk data reprocessed and put in delivery 20101206

comment:25 Changed 8 years ago by emca

Hawk data has been checked and is ready for re-delivery

comment:26 Changed 8 years ago by anch

  • Description modified (diff)

Hawk redelivered 20/12/2010

comment:27 Changed 7 years ago by jaho

Hyperspectral SCT's

line eagle hawk
2 0.02 0.00
3 0.01 0.01
4 0.04 -0.01
5 0.05 0.00
6 0.01 0.04
7 0.07 -0.01
8 -0.03 0.01
9 0.03 0.00
10 0.05 0.00
11 0.01 0.01
12 0.02 -0.01
13 0.07 0.02
14 0.02 -0.04
15 -0.01 0.08
16 0.05 -0.05
17 -0.03 1.1

comment:28 Changed 6 years ago by anhi

Reprocessing delivery check.

Logsheet needs renaming to GB10_07-2010_063-Upavon.pdf (_ to - in front of Upavon)

Everything else looks good.

comment:29 Changed 6 years ago by anhi

Reprocessing hyperspectral delivered to Kate Welham on 04/10/12 on disk 4

comment:30 Changed 6 years ago by anhi

Reprocessing hyperspectral uploaded to ftp site on 04/10/12

comment:31 Changed 5 years ago by knpa

  • Owner changed from anch to knpa
  • Status changed from accepted to assigned
  • Summary changed from GB10/07, flight day 63/2010, Upavon to GB10/07, flight day 063/2010, Upavon

Beginning archiving.

comment:32 Changed 5 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Backed-up by NEODC.

Note: See TracTickets for help on using tickets.